OpenStudy (anonymous):

w squared+3w-4 can you help me on this

OpenStudy (anonymous):

is is to be factorised or what ?????

OpenStudy (turingtest):

\[w^2+3w-4=0\]I'm guessing, so\[(w+4)(w-1)=0\rightarrow w=\left\{ -4,1 \right\}\]